# Cartesian atomic cluster expansion for machine learning interatomic potentials

### v2

@@ Abstract
Machine learning interatomic potentials are revolutionizing large-scale, accurate atomistic modelling in material science and chemistry. Many potentials use atomic cluster expansion or equivariant message passing frameworks. Such frameworks typically use spherical harmonics as angular basis functions, and then use Clebsch-Gordan contraction to maintain rotational symmetry, which may introduce redundancies in representations and computational overhead. We propose an alternative: a Cartesian-coordinates-based atomic density expansion. This approach provides a complete set of polynormially indepedent features of atomic environments while maintaining interaction body orders. Additionally, we integrate low-dimensional embeddings of various chemical elements and inter-atomic message passing. The resulting potential, named Cartesian Atomic Cluster Expansion (CACE), exhibits good accuracy, stability, and generalizability. We validate its performance in diverse systems, including bulk water, small molecules, and 25-element high-entropy alloys.
I Introduction
@@
Notably, atomic cluster expansion (ACE) Drautz (2019) represents atomic environments based on a body-ordered expansion, utilizing a complete basis set of spherical harmonics and radial components. ACE can be viewed as a general framework of many other atomic environment representations, such as Atom Centered Symmetry Functions (ACSF) Behler and Parrinello (2007), Smooth Overlap of Atomic Positions (SOAP) descriptor Bartók et al. (2010), Moment Tensor Potentials (MTPs) Shapeev (2016), and bispectrums Bartók et al. (2013).
@@
Another important approach for learning atomic interactions uses message passing neural networks (MPNNs). They represent structures as graphs with atoms as nodes, and apply message passing operations to learn atomic environment representations. Earlier MPNN models, such as SchNet Schütt et al. (2017), PhysNet Unke and Meuwly (2019), SphereNet Liu et al. (2022) and GemNet Gasteiger et al. (2021), use internal features that are invariant under rotations. In models such as NewtonNet Haghighatlari et al. (2022), EGNN Satorras et al. (2021), and PaiNN Schütt et al. (2021), vector features built using relative atomic positions are also used on top of invariant features. In later equivariant MPNNs, such as Cormorant Anderson et al. (2019), NequIP Batzner et al. (2022), MACE Batatia et al. (2022a) exploit internal features based on irreducible representations of the E(3) symmetry group Geiger and Smidt (2022); Goodman and Wallach (2000).
@@
Under the hood, in both ACE and E(3) equivariant MPNNs, the key process is the following: Equivariant features or messages of body order $\nu$ are encoded in spherical harmonics with degrees $l$. The invariant features are subsequently created by contracting the equivariant ones via the Clebsch-Gordan coefficients. These invariants are used to predict energies or other invariant physical quantities. Although the set of spherical harmonics forms a natural basis for the irreducible representations of the SO(3) group and thus is useful in operations involving rotational symmetries, it has drawbacks: the transformation of atomic coordinates to a spherical coordinate system can be an overhead although using Legendre polynomials alleviates this problem Drautz (2020), and the Clebsch-Gordan contraction is non-intuitive and cumbersome to use. As the degree $l$ and the body order $\nu$ increases, the matrix of Clebsch-Gordan coupling coefficients becomes very large ($\mathcal{O}(l+1)^{2\nu}$) but extremely sparse, and the number of the equivariant features increases exponentially when explicitly evaluated Darby et al. (2023); Nigam et al. (2020).
@@
Here, we propose an alternative method that performs the expansion of atomic density as well as symmetrization of orientation-dependent features in Cartesian coordinates directly, circumventing spherical harmonics. The Cartesian atomic cluster expansion (CACE) has the same mathematical foundation as the ACE framework, so it also adopts the nice properties of ACE – body order and completeness in describing atomic environments Dusson et al. (2022). The CACE invariant features are low-dimensional and can be computed independently and efficiently. By further incorporating a low-dimensional embedding for chemical elements and message passing between neighboring atoms, we develop the CACE potential. We benchmark CACE on diverse systems, with emphasis on stability and extrapolation.
II Methods
Atomic graph
@@ II Methods
As illustrated in Fig. 1, each atomic structure is represented as a graph, with atoms as nodes, and directed edges that point from pairs of sender atoms to receiver atoms that are within a cutoff radius of $r_{\mathrm{cut}}$. The state of an atom $i$ is denoted by the tuple $\sigma_{i}=(\textbf{r}_{i},\boldsymbol{\theta}_{i},\textbf{h}_{i})$: $\textbf{r}_{i}$ is its Cartesian position, $\boldsymbol{\theta}_{i}$ encodes its chemical element $\alpha$, and $\textbf{h}_{i}$ is the learnable features. Later we will use the superscript, e.g. $\sigma_{i}^{(t)}$, to denote the atomic state in the message passing layer $t$, but for now we omit the layer-dependence indices for simplicity.
@@
For both the sender node and the receiver node of an edge, the elements $\alpha_{i}$ are first mapped to vectors of lengths equal to the total number of elements $N_{\mathrm{element}}$ via one-hot embedding, and then the one-hot vectors are multiplied with a learnable weight matrix of size $N_{\mathrm{embedding}}\times N_{\mathrm{element}}$ outputting a learnable embedding $\boldsymbol{\theta}_{i}$ for each atom $i$ of length $N_{\mathrm{embedding}}$. Typically, $N_{\mathrm{embedding}}$ is selected to be a small number between 1 and 4. This continuous element embedding eliminates the unfavorable scaling of using discrete chemical element types when representing atomic environments, and also allows alchemical learning.
Edge basis
The edge basis, shown in Fig. 1b, is used to describe the spatial arrangement of an atom $j$ around the atom $i$ ($\vec{\mathbf{r}}_{ji}\equiv\mathbf{r}_{j}-\mathbf{r}_{i}$, $r_{ji}=|\vec{\mathbf{r}}_{ji}|$ and $\hat{\mathbf{r}}_{ji}=\vec{\mathbf{r}}_{ji}/r_{ji}$), and is formed by the product of a radial basis $R$, an angular basis $L$, and an edge-type basis $T$:
@@
In CACE, the type of edge $T$ depends on the states of the two nodes it connects. In the initial layer, the edge states are only determined by the chemical elements of the two atoms $i$ and $j$. The edge is encoded using the flattened tensor product of the two node embedding vectors, $T=\boldsymbol{\theta}_{i}\otimes\boldsymbol{\theta}_{j}$, resulting in an edge feature of length $c=N_{\mathrm{embedding}}^{2}$. In a later message passing layer $t$, the edge features can depend on the node hidden features $\mathbf{h}_{i}^{(t)}$. Such edge encoding formed by the tensor product can be interpreted as the attention mechanism Vaswani et al. (2017) between the two atoms, with the sender and the receiver node embedding vectors being the query and the key.
@@
where $\mathbf{l}\equiv(l_{x},l_{y},l_{z})$ are the angular momentum numbers satisfying $l_{x}+l_{y}+l_{z}=l$, with $l$ being the total angular momentum. Such angular basis is just spherical harmonic functions in Cartesian basis. Indeed, one can easily transform the function of a certain angular momentum number $l$ in Cartesian and spherical harmonic bases by a simple matrix multiplication Altmann (1957).
$R_{n,cl}(r_{ji})$ is a set of radial basis consisting of $n$ functions, coupled with the total angular momentum $l$ and the edge channel $c$. This means that, for each $l$ and $c$ combination, there is a different set of $n$ radial functions. In CACE, the raw radial embedding block uses a set of trainable Bessel functions with different frequencies Schütt et al. (2021) multiplied by a smooth cutoff function. The output is then mixed using a linear transformation to form $R_{n,cl}$ for each $l$ and $k$.
Atom-centered basis ($A$-basis)
As shown in Fig. 1c, the atom-centered representation is made by summing over all the edges of a node,
@@
Symmetrized basis ($B$-basis)
The orientation-dependent $A$ features are symmetrized to get the invariant $B$ features of different body orders $\nu$ as sketched in Fig. 1c. For $l=0$, $A$ is already rotationally invariant, so the $\nu=1$ features $B^{(1)}_{i,cn0}=A_{i,cn\textbf{0}}$.
For $\nu\geq 2$, the general rule for constructing the symmetrized basis is illustrated in Fig. 1i. The key idea is that pairs of angular features need to have a shared factor of $\mathbf{l}$, when summed over to form invariants. The terms that only differ from others by permutations of the $l$-combinations are not considered. Importantly, if a combination contains any elements satisfying $\mathbf{l}=\mathbf{0}$ or $\mathbf{l}=\mathbf{l^{\prime}}$, the resulting invariant can be constructed using the product of lower-order invariants, and can thus be eliminated for redundancy. The CACE framework thus makes polynomially independent invariants, which are much more compact than ACE Drautz (2019) or MTPs Shapeev (2016), where the basis functions span the space of all invariants including linearly dependent terms Dusson et al. (2022). Other radial or edge terms that do not depend on $\mathbf{l}$ or only depend on the total angular momentum $l$ can be multiplied on top without breaking the rotational symmetry.
More specifically, for $\nu=2$ Zhang et al. (2021a),
$B_{i,cnl_{1}}^{(2)}=\sum_{\mathbf{l}_{1}}\mathcal{C}(\mathbf{l}_{1})A_{i,cn\mathbf{l}_{1}}^{2},$ (4)
summed over all $\mathbf{l}_{1}$ with $l_{1x}+l_{1y}+l_{1z}=l_{1}$. $\mathcal{C}(\mathbf{l})=\dfrac{l!}{l_{x}!l_{y}!l_{z}!}$ is the combinatorial prefactor. Doing the expansion explicitly, one can use the multinomial formula to show that Zhang et al. (2021a)
@@
Based on the expansion, one can see that these $B^{(\nu)}$ terms are the same as the $B^{(\nu)}$ terms in the original ACE paper Drautz (2019), although many linearly or polynomially-dependent features are eliminated and the computation in the Cartesian space is much more straightforward. The $B$ basis is a complete basis of permutation and rotation invariant functions of the atomic environment Drautz (2019). In practice, one truncates the series up to some maximum values of $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$ to make the computation tractable. The dimension of the $B$ basis is $c\times n\times N_{L}$, where each factor comes from the number of edge channels, radial basis, and angular basis, respectively. Fig. 2 shows the total number of angular features $N_{L}$ up to certain $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$. The number of $N_{l}$ is compact even for $\nu_{\mathrm{max}}=5$. For most practical applications, $l_{\mathrm{max}}$ between 2 and 4 and $\nu_{\mathrm{max}}$ between 2 and 4 are sufficient for good accuracy when fitting CACE potentials.
@@
Message passing
The framework above reproduces the original ACE, albeit formulated entirely in the Cartesian space. In ACE, the atomic energy of each atom is determined by the neighboring atoms within its cutoff radius $r_{\mathrm{cut}}$. On top of that, one can incorporate message passing, which iteratively propagates information along the edges, thereby communicating the local information and effectively enlarging the perceptive field of each atom. A systematic discussion on the design choices of message passing can be found in Refs. Batatia et al. (2022b); Bochkarev et al. (2024).
@@
$M_{i,cn\mathbf{l}}^{(t)}=\sum_{j\in\mathcal{N}(i)}m_{1,j\rightarrow i,kn\mathbf{l}}^{(t)}+m_{2,j\rightarrow i,kn\mathbf{l}}^{(t)}$ (10)
@@
and for $G$ we use a simple linear function. After obtaining the updated $A^{(t+1)}$, we again take the symmetrized $B$-basis as described before. It is easy to verify that these symmetrized representations are only dependent on the scalar values of the atomic distances and angles, and thus rotationally invariant. Note that one can make various design choices for $F$, $H$ and $G$ functions: they can be dependent on $c$, $n$ and $l$, without breaking the invariance at the symmetrization stage.
@@
Readout and output
After $T$ MP layers, all the resulting $B$ features at each layer are concatenated. These invariant features are compact; the number of the features can be computed as $T\times c\times n\times N_{L}$. Then, a multilayer perceptron (MLP) maps these invariant features to the target of the atomic energy of each atom $i$,
@@
In practice, for the readout function we use the sum of a linear NN and a multilayer NN, as the former preserves the body-ordered contributions and the latter considers the higher-ordered terms from the truncated series.
@@
III Applications
III.1 Bulk water
As an example application on complex fluids, we demonstrate CACE on a dataset of 1,593 liquid water configurations, each of 64 molecules Cheng et al. (2019). The dataset was computed at the revPBE0-D3 level of density functional theory (DFT). 90% of the data are used in the training. We used a cutoff $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions and $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and no message passing ($T=0$) or one message passing layer ($T=1$). The fitted CACE models have relatively few parameters: 24,572 trainable parameters for the CACE water model with $T=0$, and 69,320 trainable parameters for the model with $T=1$. The root mean square errors (RMSE) of the energy per water molecule and forces on the validation set are presented in Table 1.
[…]
@@ III.1 Bulk water
[…]
Refer to captionaRefer to caption Refer to captionb

### v3

@@ Abstract
Machine learning interatomic potentials are revolutionizing large-scale, accurate atomistic modelling in material science and chemistry. Many potentials use atomic cluster expansion or equivariant message passing frameworks. Such frameworks typically use spherical harmonics as angular basis functions, followed by Clebsch-Gordan contraction to maintain rotational symmetry. We propose a mathematically equivalent and simple alternative that performs all operations in the Cartesian coordinates . This approach provides a complete set of polynormially independent features of atomic environments while maintaining interaction body orders. Additionally, we integrate low-dimensional embeddings of various chemical elements, trainable radial channel coupling, and inter-atomic message passing. The resulting potential, named Cartesian Atomic Cluster Expansion (CACE), exhibits good accuracy, stability, and generalizability. We validate its performance in diverse systems, including bulk water, small molecules, and 25-element high-entropy alloys.
@@
In particular, atomic cluster expansion (ACE) Drautz (2019) represents atomic environments based on a body-order expansion, using a complete set of bases of spherical harmonics and radial components. ACE can be viewed as a general framework of many other representations of the atomic environment, such as the Atom Centered Symmetry Functions (ACSF) Behler and Parrinello (2007), Smooth Overlap of Atomic Positions (SOAP) descriptor Bartók et al. (2010), Moment Tensor Potentials (MTPs) Shapeev (2016), and bispectrums Bartók et al. (2013).
@@
Another important approach to learning atomic interactions uses message passing neural networks (MPNNs). They represent structures as graphs with atoms as nodes, and apply message passing operations to learn atomic environment representations. Earlier MPNN models, such as SchNet Schütt et al. (2017), PhysNet Unke and Meuwly (2019), SphereNet Liu et al. (2022) and GemNet Gasteiger et al. (2021), use internal features that are invariant under rotations. In models such as NewtonNet Haghighatlari et al. (2022), EGNN Satorras et al. (2021), and PaiNN Schütt et al. (2021), vector features built using relative atomic positions are also used on top of invariant features. In later equivariant MPNNs, such as Cormorant Anderson et al. (2019), NequIP Batzner et al. (2022), MACE Batatia et al. (2022a) exploit internal features based on irreducible representations of the E(3) symmetry group Geiger and Smidt (2022); Goodman and Wallach (2000). E(3) equivariant MPNNs were able to achieve an unprecedented accuracy compared to the previous invariant architectures Batzner et al. (2022); Batatia et al. (2022a).
@@
Under the hood, in both ACE and E(3) equivariant MPNNs, the key process is the following: Equivariant features or messages of body order $\nu$ are encoded in spherical harmonics with degrees $l$. The invariant features are subsequently created by contracting the equivariant ones via the Clebsch-Gordan coefficients. These invariants are used to predict energies or other invariant physical quantities. The use of spherical harmonics is motivated by that they forms a natural basis for the irreducible representations of the SO(3) group and thus is useful in operations involving rotational symmetries. Since the original ACE work Drautz (2019), various improvements have been made, e.g., efficient implementations of spherical harmonics that are evaluated in Cartesian coordinates are made Sloan (2013); Drautz (2020); Bigi et al. (2023), absorbing Clebsch-Gordan matrix in expansion coefficients during the evaluation stage Lysogorskiy et al. (2021), recursive Clebsch-Gordan coupling Batatia et al. (2022a) or using Gaunt coefficients Luo et al. (2024). Moreover, the original formulation has a steep scaling as the number of elements increases, and tensor-reduced representations have been developed to combat this Darby et al. (2023). As the degree $l$ and the body order $\nu$ increases, many invariant features are linearly dependent or polynomially depedent, and schemes to eliminate these have been developed Dusson et al. (2022); Nigam et al. (2020); Goff et al. (2024).
On the other hand, MTP Shapeev (2016) can be considered as ACE formulated in the Cartesian space. MTP uses the so-called moment tensors, and then contracts them to form rotationally invariant basis functions Shapeev (2016). The MTP contraction yields a spanning set, but contains many linearly dependent basis functions Shapeev (2016); Dusson et al. (2022). In a similar vain, proper orthogonal descriptors were introduced as compact orthogonal basis functions formulated in the Cartesian space Nguyen (2023).
@@
Here, we propose an alternative method that performs the expansion of atomic density as well as the contraction to get polynormially indepedent invariant features in Cartesian coordinates directly, circumventing spherical harmonics. The Cartesian atomic cluster expansion (CACE) has the same mathematical foundation as the ACE framework, so it also adopts the nice properties of ACE: body order and completeness in the description of atomic environments Dusson et al. (2022). The CACE invariant features are low-dimensional and can be computed independently and efficiently. By further incorporating a low-dimensional embedding for chemical elements, optimized radial channel coupling, and message passing between neighboring atoms, we develop the CACE potential. We benchmark CACE on diverse systems, with an emphasis on stability and extrapolation.
I Results
I.1 The CACE architecture
@@
Atomic graph As illustrated in Fig. 1, each atomic structure is represented as a graph, with atoms as nodes, and directed edges that point from pairs of sender atoms to receiver atoms that are within a cutoff radius of $r_{\mathrm{cut}}$. The state of an atom $i$ is denoted by the tuple $\sigma_{i}=(\textbf{r}_{i},\boldsymbol{\theta}_{i},\textbf{h}_{i})$: $\textbf{r}_{i}$ is its Cartesian position, $\boldsymbol{\theta}_{i}$ encodes its chemical element $\alpha$, and $\textbf{h}_{i}$ is the learnable features. Later we will use the superscript, e.g. $\sigma_{i}^{(t)}$, to denote the atomic state in the message passing layer $t$, but for now we omit the layer-dependence indices for simplicity.
@@
For both the sender node and the receiver node of an edge, elements $\alpha_{i}$ are first assigned to vectors of lengths equal to the total number of elements $N_{\mathrm{element}}$ through a one-hot embedding, and then the one-hot vectors are multiplied with a learnable weight matrix of size $N_{\mathrm{embedding}}\times N_{\mathrm{element}}$ that outputs a learnable embedding $\boldsymbol{\theta}_{i}$ for each atom $i$ of length $N_{\mathrm{embedding}}$. Typically, $N_{\mathrm{embedding}}$ is selected to be a small number between 1 and 4. This continuous element embedding eliminates the unfavorable scaling of using discrete chemical element types when representing atomic environments, and also allows alchemical learning.
Edge basis The edge basis, shown in Fig. 1b, is used to describe the spatial arrangement of an atom $j$ around the atom $i$ ($\vec{\mathbf{r}}_{ji}\equiv\mathbf{r}_{j}-\mathbf{r}_{i}$, $r_{ji}=|\vec{\mathbf{r}}_{ji}|$ and $\hat{\mathbf{r}}_{ji}=\vec{\mathbf{r}}_{ji}/r_{ji}$), and is formed by the product of a radial basis $R$, an angular basis $L$, and an edge-type basis $T$:
@@
In CACE, the type of edge $T$ depends on the states of the two nodes it connects. In the initial layer, the edge states are only determined by the chemical elements of the two atoms $i$ and $j$. The edge is encoded using the flattened tensor product of the two node embedding vectors, $T=\boldsymbol{\theta}_{i}\otimes\boldsymbol{\theta}_{j}$, resulting in an edge feature of length $c=N_{\mathrm{embedding}}^{2}$. In a later message passing layer $t$, the edge features can depend on the node hidden features $\mathbf{h}_{i}^{(t)}$. This edge encoding formed by the tensor product can be interpreted as the attention mechanism Vaswani et al. (2017) between the two atoms, with the embedding vectors of the sender and the receiver nodes the query and the key.
@@
where $\mathbf{l}\equiv(l_{x},l_{y},l_{z})$ are the angular momentum numbers satisfying $l_{x}+l_{y}+l_{z}=l$, with $l$ being the total angular momentum. Such angular basis is just spherical harmonic functions in Cartesian basis. Indeed, one can easily transform the function of a certain angular momentum number $l$ in Cartesian and spherical harmonic bases by a simple matrix multiplication Altmann (1957). As examples of operations in the Cartesian angular basis, one can add two vectors such as
$L_{\mathbf{l}_{3}}(\hat{\mathbf{r}}_{ji}+\hat{\mathbf{r}}_{ik})=\sum_{\mathbf{l}_{1}+\mathbf{l}_{2}=\mathbf{l}_{3}}\mathcal{P}(\mathbf{l}_{1},\mathbf{l}_{2})L_{\mathbf{l}_{1}}(\hat{\mathbf{r}}_{ji})L_{\mathbf{l}_{2}}(\hat{\mathbf{r}}_{ik})$ (3)
where the prefactor $\mathcal{P}(\mathbf{l}_{1},\mathbf{l}_{2})=\dfrac{l_{3x}!}{l_{1x}!l_{2x}!}\dfrac{l_{3y}!}{l_{1y}!l_{2y}!}\dfrac{l_{3z}!}{l_{1z}!l_{2z}!}$ with $\mathbf{l}_{1}+\mathbf{l}_{2}=\mathbf{l}_{3}$. One can obtain vector dot products such as
$(\hat{\mathbf{r}}_{ji}\cdot\hat{\mathbf{r}}_{ki})^{l}=\sum_{\mathbf{l}}\mathcal{C}(\mathbf{l})L_{\mathbf{l}}(\hat{\mathbf{r}}_{ji})L_{\mathbf{l}}(\hat{\mathbf{r}}_{ki}),$ (4)
where $\mathcal{C}(\mathbf{l})=\dfrac{l!}{l_{x}!l_{y}!l_{z}!}$ is the combinatorial coefficient, and the sum is over all $\mathbf{l}$ with $l_{x}+l_{y}+l_{z}=l$. One can prove the relationships in Eqns. (3) and  (4) above using the multinomial formula.
$R_{n,cl}(r_{ji})$ is a set of radial basis consisting of $n$ functions, coupled with the total angular momentum $l$ and the edge channel $c$. CACE first uses a set of raw radial basis $R_{\tilde{n}}(r_{ji})$ consisting of $\tilde{n}$ functions, which can be a set of trainable Bessel functions with different frequencies Schütt et al. (2021) multiplied by a smooth cutoff function. The raw radial basis is mixed using a linear transformation to form $R_{n,cl}$ for each $c$ and $l$,
$R_{n,cl}(r_{ji})=\sum_{\tilde{n}}R_{\tilde{n}}(r_{ji})W_{\tilde{n}n,cl}$ (5)
where $W_{\tilde{n}n,cl}$ is a $\tilde{n}\times n$-sized matrix indexed by $c$ and $l$. This means that, for each $l$ and $c$ combination, there is a different set of $n$ radial functions. In practice, the mixing is performed on the atom-centered basis for efficiency, as described below.
Atom-centered basis ($A$-basis) As shown in Fig. 1c, the atom-centered representation is made by summing over all the edges of a node,
@@
Symmetrized basis ($B$-basis) The orientation-dependent $A$ features are symmetrized to get the invariant $B$ features of different body orders $\nu$ as sketched in Fig. 1c. For $l=0$, $A$ is already rotationally invariant, so the $\nu=1$ features $B^{(1)}_{i,cn0}=A_{i,cn\textbf{0}}$. For $\nu=2$ Zhang et al. (2021a),
$B_{i,cnl_{1}}^{(2)}=\sum_{\mathbf{l}_{1}}\mathcal{C}(\mathbf{l}_{1})A_{i,cn\mathbf{l}_{1}}^{2}.$ (7)
Doing the expansion explicitly and using Eqn. (3), one can show that Zhang et al. (2021a)
@@
The general rule for constructing the symmetrized basis is illustrated in Fig. 1i. The key idea is that pairs of angular features $L$ need to have a shared factor of $\mathbf{l}$, when summed over to form invariants. To eliminate linearly dependent terms, the combination of the angular features that only differ from others by permutations are not considered. Moreover, the shared factor $\mathbf{l}$ between any pairs of $L$ needs to be greater than zero, because if any $\mathbf{l}=\mathbf{0}$ the resulting invariant can be constructed using the product of lower-order invariants and can thus be eliminated for redundancy. In other words, a valid combination illustrated in Fig. 1i needs to correspond to one connected graph, rather than two disconnect graphs of lower body orders. For instance, terms such as $B_{i,cn012}^{(4)}=B_{i,cn0}^{(1)}B_{i,cn12}^{(3)}$, $B_{i,cn102}^{(4)}=B_{i,cn1}^{(2)}B_{i,cn2}^{(2)}$ are eliminated. The CACE framework thus makes polynomially independent invariants. Other radial or edge terms that do not depend on $\mathbf{l}$ or only depend on the total angular momentum $l$ can be multiplied on top without breaking the rotational symmetry.
Based on the expansions, one can see that these CACE $B^{(\nu)}$ terms are similar to the $B^{(\nu)}$ terms in the original ACE paper Drautz (2019) or the invariant basis functions in MTP Shapeev (2016) , with two differences: First, many linearly or polynomially-dependent features are elimiated during the symmetrization stage in CACE, resulting in a much more compact basis set than the original ACE  Drautz (2019) or MTPs Shapeev (2016), where the basis functions span the space of all invariants including linearly dependent terms Dusson et al. (2022). Second, since each radial channel $R_{n,cl}$ contains all raw radial channels with trainable weights (Eqn. (5)), it is not necessary to couple different radial channels together when forming the basis $B$. This means that the number of radial channels of $B$ stays $n$ rather than going up as $n^{\nu}$ as the body order increases.
The $B$ basis is a complete basis of permutation and rotation invariant functions of the atomic environment Drautz (2019). In practice, one truncates the series up to some maximum values of $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$ to make the computation tractable. The dimension of the $B$ basis is $c\times n\times N_{L}$, where each factor comes from the number of edge channels, radial basis, and angular basis, respectively. Fig. 2 shows the total number of angular features $N_{L}$ up to certain $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$. The number of $N_{l}$ is compact even for $\nu_{\mathrm{max}}=5$. For most practical applications, $l_{\mathrm{max}}$ between 2 and 4 and $\nu_{\mathrm{max}}$ between 2 and 4 are sufficient for good accuracy when fitting CACE potentials.
@@
Message passing The framework above reproduces the original ACE, although formulated entirely in Cartesian space. In ACE, the atomic energy of each atom is determined by neighboring atoms within its cutoff radius $r_{\mathrm{cut}}$. On top of that, one can incorporate message passing, which iteratively propagates information along the edges, thereby communicating the local information and effectively enlarging the perceptive field of each atom. A systematic discussion on the design choices of message passing can be found in Refs. Batatia et al. (2022b); Bochkarev et al. (2024).
@@
$M_{i,cn\mathbf{l}}^{(t)}=\sum_{j\in\mathcal{N}(i)}m_{1,j\rightarrow i,cn\mathbf{l}}^{(t)}+m_{2,j\rightarrow i,cn\mathbf{l}}^{(t)}$ (13)
@@
and for $G$ we use a simple linear function. After obtaining the updated $A^{(t+1)}$, we again take the symmetrized $B$-basis as previously described. It is easy to verify that these symmetrized representations are dependent only on the scalar values of the atomic distances and angles and thus rotationally invariant. Note that one can make various design choices for the $F$, $H$ and $G$ functions: they can be dependent on $c$, $n$, and $l$, without breaking the invariance at the symmetrization stage.
@@
Readout and output After $T$ MP layers, all the resulting $B$ features at each layer are concatenated. These invariant features are compact; the number of the features can be computed as $T\times c\times n\times N_{L}$. Then, a multilayer perceptron (MLP) maps these invariant features to the target of the atomic energy of each atom $i$,
@@
In practice, for the readout function, we use the sum of a linear NN and a multilayer NN, as the former preserves the body-ordered contributions and the latter considers the higher-ordered terms from the truncated series.
@@
Table 1: Root mean square errors (RMSE) for energy (E) per water molecule (in meV/H2O) and force (in meVÅ-1) on a liquid water dataset from Ref. Cheng et al. (2019). The cutoffs $r_{\mathrm{cut}}$ of the atomic environment and the numbers of message passing layers $T$ are listed. The NequIP and DeepMD models are trained by us, and the others are from the references.
@@
I.2 Bulk water
As an example application on complex fluids, we demonstrate CACE on a dataset of 1,593 liquid water configurations Cheng et al. (2019). The details of the training are in the Methods section. The root mean square errors (RMSE) of the energy per water molecule and forces on the validation set for the CACE model and other MLIPs are presented in Table 1. MLIPs based on three-body and three-body descriptors without message passing, including BPNN based on ACSF Behler and Parrinello (2007), embedded atom neural network (EANN) Zhang et al. (2021b), and DeepMD have similar errors, and are amongst the highest in this comparison. Linear ACE uses higher-body order features, which can be directly compared to the CACE model trained with the same cutoff and without message passing ($r_{\mathrm{cut}}=5.5$, $T=0$), and the latter has lower error. The REANN potential Zhang et al. (2021a) with three message passing layers, the equivariant message passing model NequIP Batzner et al. (2022), the high body order MACE Batatia et al. (2022a) model, and the CACE model with $T=1$ have the lowest errors.
For running molecular dynamics (MD) simulations using MLIPs, besides accuracy, a crucial requirement is stability, the possibility to run long enough trajectories with the system remaining in physically reasonable states and the simulation not exploding, even at conditions with few or no training data. To assess the simulation stability, we performed independent simulations of 300 ps in length for water at 1 gmL-1 at 300 K, 500 K, 1000 K, 1500 K, and 2000 K using the CACE $T=1$ model. The simulation cell contains 512 water molecules. The time step was 1 femtosecond and the Nosé-Hoover thermostat was used. The upper panel of Fig. 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF). At 300 K, the computed O-O RDF is in excellent agreement with experiment from X-ray diffraction measurements Skinner et al. (2014). Note that the nuclear quantum effects have a slight destructuring effect in the O-O RDF, but the effect is quite small Cheng et al. (2019). The mean squared displacements (MSD) from these simulations are shown on the lower panel of Fig. 3. The diffusivity $D$, obtained by fitting a linear function to the MSD, at 300 K agrees well with the DFT MD result ($D_{\mathrm{DFT}}=2.67\pm 0.10\mathrm{\AA}^{2}$ps-1) from Ref. Marsalek and Markland (2017). At high temperatures up to 2000 K, the MD simulations remained stable, which demonstrates the superior stability of the CACE potential. Note that we do not expect the result at very high temperatures to be physically predictive, and this example is just to illustrate the stability of CACE. Such stability is not only necessary for converging statistical properties, but also allows the refinement of the potential: to iteratively improve the accuracy of the potential under certain conditions, one can perform active learning which collects snapshots from corresponding MD trajectories and use these snapshots for re-training.